# Grassroots Platforms with Atomic Transactions: Social Networks, Cryptocurrencies, and Democratic Federations

### v3

@@ Abstract
Here, we aim to provide a more suitable formal foundation for grassroots platforms. To do so, we enhance the notion of a distributed transition system to include atomic transactions and revisit the notion of grassroots platforms within this new foundation. We present crisp specifications of key grassroots platforms using atomic transactions: befriending and defriending for grassroots social networks, coin swaps for grassroots cryptocurrencies, and communities forming, joining, and leaving a federation for grassroots democratic federations. We prove a general theorem that a platform specified by atomic transactions that are so-called interactive is grassroots; show that the atomic transactions used to specify all three platforms are interactive; and conclude that the platforms thus specified are indeed grassroots. We thus provide a crisp mathematical foundation for grassroots platforms and a solid and clear starting point from which their implementation can commence.
@@ 1 Introduction
1. Befriend: $c^{\prime}_{p}:=c_{p}\cup\{q\}$, $c^{\prime}_{q}:=c_{q}\cup\{p\}$, provided $q\notin c_{p}$ and $p\notin c_{q}$ 2. Defreind: $c^{\prime}_{p}:=c_{p}\setminus\{q\}$, $c^{\prime}_{q}:=c_{q}\setminus\{p\}$, provided $q\in c_{p}$ and $p\in c_{q}$
@@
Grassroots Cryptocurrencies. Grassroots cryptocurrencies [45, 29] can provide a foundation for an egalitarian digital economy, where people price their goods and services in terms of their own personally-minted grassroots coins, and liquidity is achieved through the creation of mutual credit lines among persons (natural and legal) via the swap of personal coins. Grassroots coins issued by different persons form an integrated digital economy via the sole rule in the grassroots cryptocurrencies protocol:
Coin Redemption: $p$ can redeem any $q$-coin it holds against any coin held by $q$, one-for-one.
@@ 3 Grassroots Platforms with Atomic Transactions
3.1 Grassroots Social Network via Befriending and Defriending
@@
    Defreind: $c^{\prime}_{p}:=c_{p}\setminus\{q\}$, $c^{\prime}_{q}:=c_{q}\setminus\{p\}$, provided $q\in c_{p}$ and $p\in c_{q}$
@@
The proof is by induction on the length of the run $r=c0,c_{1},\ldots,c_{n}$. Assume $|r|=1$, namely $r$ consists of the initial configuration $c0$. Then all local states are empty, satisfying the Lemma. Assume the lemma holds for runs of length $|r|=n$, that $c_{n}$ satisfies the Lemma, and consider the transition $c_{n}\rightarrow c_{n+1}$. It can be either Befriend or Defriend; both satisfy for every $p\in P$ the condition $p\in c_{q}$ iff $q\in c_{p}$ for $c_{n+1}$ if $c_{n}$ does.
@@ 3.3 Grassroots Democratic Federation via Joining and Leaving
This completes the transactions-based specification of the three key grassroots platforms. It does not distinguish between transactions carried out by mutual consent (befriend, open a credit line, join), and transactions that are obligatory once requested (defriend, redeem, leave). The distinction can be added at this level of abstraction, with a liveness requirement that once a request to carry out an obligatory transaction is made, it is eventually carried out. Alternatively, it can be added at the unary implementation level; this may be more natural as atomic transactions would anyway be realised by one agent issuing a request/offer and the other(s) consenting to it [44, 29].
@@ 5 Implementation
For now, we note that, worst case, a friendship offer by $p$ accepted by $q$, or an offer by community $p$ to join community $q$, would remain in limbo. If it is committed by $p$ at some later point, which is not convenient to $q$, then $q$ can promptly defriend $q$ or remove $q$, as the case may be, with little or not harm done. In case of grassroots cryptocurrencies, a swap transaction in limbo may tie coins offered by $q$, which may or may not be harmful to $q$ (not harmful if these are $q$-coins, which $q$ may mint as it pleases; or $p$-coins that $q$ tries to redeem, and if $p$ is non-responsive it might indicate that $p$-coins are not worth much anyhow).

### v4

@@ Abstract
Here, we aim to provide a more suitable formal foundation for grassroots platforms. To do so, we enhance the notion of a distributed transition system to include atomic transactions and revisit the notion of grassroots platforms within this new foundation. We present crisp specifications of key grassroots platforms using atomic transactions: befriending and unfriending for grassroots social networks, coin swaps for grassroots cryptocurrencies, and communities forming, joining, and leaving a federation for grassroots democratic federations. We prove a general theorem that a platform specified by atomic transactions that are so-called interactive is grassroots; show that the atomic transactions used to specify all three platforms are interactive; and conclude that the platforms thus specified are indeed grassroots. We thus provide a crisp mathematical foundation for grassroots platforms and a solid and clear starting point from which their implementation can commence.
@@ 1 Introduction
1. Befriend: $c^{\prime}_{p}:=c_{p}\cup\{q\}$, $c^{\prime}_{q}:=c_{q}\cup\{p\}$, provided $q\notin c_{p}$ and $p\notin c_{q}$ 2. Unfriend: $c^{\prime}_{p}:=c_{p}\setminus\{q\}$, $c^{\prime}_{q}:=c_{q}\setminus\{p\}$, provided $q\in c_{p}$ and $p\in c_{q}$
@@
Grassroots Cryptocurrencies. Grassroots cryptocurrencies [45, 29] can provide a foundation for an egalitarian digital economy, where people price their goods and services in terms of their own personally-minted grassroots coins, and liquidity is achieved through the creation of mutual credit lines among persons (natural and legal) via the swap of personal coins. Grassroots coins issued by different persons form an integrated digital economy via the sole rule in the grassroots cryptocurrencies protocol: [^1] This redemption rule is weaker than the one presented in previous work [45]. The reason is that as long as $q$ holds $p$-coins it can dodge a redemption claim by $p$ by issuing a counter-claim, until the mutual credit line between $p$ and $q$ dries up. Following that, $p$ can force redemption only if, to begin with, $p$ had a positive credit balance with $q$, namely held more $q$-coins than the number of $p$-coins held by $q$. This weaker redemption obligation eschews such ‘redemption wars’: It allows $p$ to redeem $q$-coins it holds against non-$p$-coins held by $q$ when their credit line ceases to be mutual, while of course not excluding voluntary swaps by mutual consent at any time, which we expect to be the norm.
Coin Redemption: $p$ can redeem a $q$-coin it holds against a $p$-coin held by $q$, if there is one, else against any coin held by $q$, one-for-one.
@@ 3 Grassroots Platforms with Atomic Transactions
3.1 Grassroots Social Network via Befriending and Unfriending
@@
    Unfriend: $c^{\prime}_{p}:=c_{p}\setminus\{q\}$, $c^{\prime}_{q}:=c_{q}\setminus\{p\}$, provided $q\in c_{p}$ and $p\in c_{q}$
@@
The proof is by induction on the length of the run $r=c0,c_{1},\ldots,c_{n}$. Assume $|r|=1$, namely $r$ consists of the initial configuration $c0$. Then all local states are empty, satisfying the Lemma. Assume the lemma holds for runs of length $|r|=n$, that $c_{n}$ satisfies the Lemma, and consider the transition $c_{n}\rightarrow c_{n+1}$. It can be either Befriend or Unfriend; both satisfy for every $p\in P$ the condition $p\in c_{q}$ iff $q\in c_{p}$ for $c_{n+1}$ if $c_{n}$ does.
@@ 3.3 Grassroots Democratic Federation via Joining and Leaving
This completes the transactions-based specification of the three key grassroots platforms. It does not distinguish between transactions carried out by mutual consent (befriend, open a credit line, join), and transactions that are obligatory once requested (unfriend, redeem, leave). The distinction can be added at this level of abstraction, with a liveness requirement that once a request to carry out an obligatory transaction is made, it is eventually carried out. Alternatively, it can be added at the unary implementation level; this may be more natural as atomic transactions would anyway be realised by one agent issuing a request/offer and the other(s) consenting to it [44, 29].
@@ 5 Implementation
For now, we note that, worst case, a friendship offer by $p$ accepted by $q$, or an offer by community $p$ to join community $q$, would remain in limbo. If it is committed by $p$ at some later point, which is not convenient to $q$, then $q$ can promptly unfriend $q$ or remove $q$, as the case may be, with little or not harm done. In case of grassroots cryptocurrencies, a swap transaction in limbo may tie coins offered by $q$, which may or may not be harmful to $q$ (not harmful if these are $q$-coins, which $q$ may mint as it pleases; or $p$-coins that $q$ tries to redeem, and if $p$ is non-responsive it might indicate that $p$-coins are not worth much anyhow).